A certain health magazine stated that 14% of men said they used exercise to reduce stress. At a = 0.05, test the claim wherein, from a random sample of 100 men who were selected, 10 said that they used exercise to relieve stress.


1
Expert's answer
2021-06-07T03:47:06-0400

"x=10 \\

n=100 \\

\hat{p}=\frac{x}{n} = \frac{10}{100}=0.1 \\

H_0: p=0.14 \\

H_1: p≠0.14"

Test statistic:

"Z = \frac{\hat{p}-p}{\sqrt{\frac{p(1-p)}{n}}} \\

= \frac{0.1-0.14}{\sqrt{\frac{0.14(1-0.14)}{100}}} \\

= -1.15"

P-value for a two tailed z-test using standard normal distribution table

P-value=0.25

As P-value is greater than the level of significance, we do not have sufficient evidence at 5% to reject null hypothesis.

We have sufficient evidence to say that 14 % of men said they used exercise to reduce stress.